OpenStudy (scarlettfarra2000):

I don't understand how to do the next part would someone please help me? Solve the equation below

OpenStudy (scarlettfarra2000):

\[z-\frac{ 3 }{ 2 }=-\frac{ 5 }{ 6 }\]

OpenStudy (scarlettfarra2000):

To solve the equation, isolate the variable. Add 3/2 to both sides of the equation \[z-\frac{ 3 }{ 2 }+\frac{ 3 }{ 2}=-\frac{ 5 }{ 6 }+\frac{ 3 }{ 2 }\]

OpenStudy (scarlettfarra2000):

Add the fractions on the right side \[z=-\frac{ 5 }{ 6 }+\frac{ 3 }{ 2 }\]

OpenStudy (aveline):

You need to change the denominator of 3/2. Multiply the top and bottom by 3

OpenStudy (mathmale):

I agree. The reason for this is that you must express all fractions with the same denominator before you can combine them.\[z=-\frac{ 5 }{ 6 }+\frac{ 3 }{ 2 }\]

OpenStudy (mathmale):

Taking the 2nd fraction, mult. its numerator and its denom. both by 3, as aveline has suggested:\[\frac{ 3 }{ 2 }*\frac{ 3 }{ 3 }\]

OpenStudy (mathmale):

Please share your result here.

OpenStudy (mathmale):

Scarlett, I don't mean to egg you on, but to continue helping you, I need to hear back from you once in a while.

OpenStudy (scarlettfarra2000):

\[\frac{ 9 }{ 6 }\] would be the answer right? And sorry @mathmale I step away also thanks you @Aveline for your help too

OpenStudy (aveline):

z=-5/6+3/2 We change 3/2 to 9/6, so... \[z=\frac{ -5 }{ 6 }+\frac{ 9 }{ 6 }\]

OpenStudy (scarlettfarra2000):

By adding the together we get \[\frac{ 14 }{ 6}\] would that be the answer? Or am I doing something wrong?

OpenStudy (mathmale):

But isn't -5+9 a small, positive number?

OpenStudy (aveline):

Just as mathmale said. Another way to write -5+9 is 9-5

OpenStudy (scarlettfarra2000):

Which would be 4

OpenStudy (mathmale):

Yes, and so what is your final value for x? Write the equation x= ... ?

OpenStudy (scarlettfarra2000):

it would stay 6 wouldn't it?

OpenStudy (mathmale):

z comes out to 4/6. Is there any way in which you could reduce this fraction?

OpenStudy (scarlettfarra2000):

\[\frac{ 4 }{ 6 }\div2 =\frac{ 2 }{ 3 }\]

OpenStudy (mathmale):

While your answer is 100% correct, the method used isn't. Have to divide BOTH the 4 and the 6 by 2:\[\frac{ 4/2 }{ 6/2 }\]

OpenStudy (aveline):

Remember, you're dividing by 2/2, not just 2. However, your answer is correct. Good job

OpenStudy (scarlettfarra2000):

okay i see now I didn't step the problem up right would I divided by 2

OpenStudy (mathmale):

Glad this conversation has led to better understanding for you. Take care, Scarlett!

OpenStudy (scarlettfarra2000):

Thanks @mathmale

OpenStudy (mathmale):

You're welcome!

OpenStudy (scarlettfarra2000):

Thanks again @Aveline

OpenStudy (aveline):

Np